OpenStudy (anonymous):

A TANGENT DRAWN TO Y=-(X^2 -4) AT THE POINT (1, 3) FORMS A RIGHT TRIANGLE WITH THE COORDINATE AXES. FIND THE AREA OF THE TRIANGLE. PLEASE HELP AND SHOW ALL STEPS!

OpenStudy (anonymous):

 The slope of this line is \[\frac{dy(1)}{dx}=-2\]The equation of this line is \[y-y_1=m(x-x_1)\], which will get you the points where the line crosses the axes, and thus the base length, and perpendicular height of the triangle.